OpenStudy (anonymous):

Which of the following best describes the solution to the system of equations below? -8x + 3y = -6 9x - 3y = 6 The system of equations has exactly one solution where x = 8 and y = 2. The system of equations has exactly one solution where x = 0 and y = -2. The system of equations has no solution. The system of equations has infinitely many solutions.

Nnesha (nnesha):

you should do elimination method bec you have one positive 3y and one negative so you can cancel out that

Nnesha (nnesha):

 what is -8x+9x=?? 3y -3y =?? -6+6 = ???